OpenStudy (anonymous):

Solve: F(20) = 100 ^ n + 1000 F(20) = 1000 * 1.1^n

OpenStudy (solomonzelman):

I don't quite understand the given statement and what has to be solved.

OpenStudy (anonymous):

Belinda wants to invest in an option that would help to increase her investment value by the greatest amount in 20 years. Will there be any significant difference in the value of Belinda's investment after 20 years if she uses option 2 over option 1?

OpenStudy (solomonzelman):

So, the formula for option 1 is, \(\color{#000000 }{ \displaystyle f(n)=100^n+1000 }\) And the formula for option 2 is, \(\color{#000000 }{ \displaystyle f(n)=1000\cdot(1.1)^n }\) in both formulas, \(\color{#000000 }{ \displaystyle n }\) - the number of years \(\color{#000000 }{ \displaystyle f(n) }\) - the total amount after n years

OpenStudy (solomonzelman):

am I correctly intepreting your data?

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

ok, cool

OpenStudy (solomonzelman):

But, when you plug in n=20, you are plugging it on the right side of the equation as well.

OpenStudy (solomonzelman):

Option 1 \(\color{#000000 }{ \displaystyle f(20)=100^{20}+1000 }\) Option 2 \(\color{#000000 }{ \displaystyle f(20)=1000\cdot(1.1)^{20} }\)

OpenStudy (anonymous):

Okay

OpenStudy (solomonzelman):

\(100^{20}=(10^{2})^{20} =10^{40}\) And that is \(1\times 10^{40}\) (in scientific notation) That is a ridiculously huge number!!

OpenStudy (solomonzelman):

Where as \(1.1^{20}\approx 6.73\)

OpenStudy (solomonzelman):

and even if you multiply this times a billion it will still be smaller than \(1\times 10^{40}\).

OpenStudy (anonymous):

okay, thank you

OpenStudy (anonymous):

f(n) = 100*n + 1000 f(n) = 1000*1.1^n

OpenStudy (solomonzelman):

Don't want to disturb, but why are you reposting these functions again?